# CARE: a Benchmark Suite for the Classification and Retrieval of Enzymes

### v1

@@ 5 Task 2: Enzyme Retrieval
Task 2, enzyme retrieval from a query reaction, tests the ability of a model to extrapolate to unseen reactions. Task 2 has not been formalized or explored in previous studies, but it is equally as important as Task 1, as it applies to a use case where a scientist or engineer is seeking to identify a previously characterized enzyme sequence that can perform a novel (unannotated) reaction (Figure 1D). Typical applications include: an environmental engineer looking for an enzyme to degrade a toxic pollutant [63], an enzyme engineer looking for an enzyme to catalyze a selective reaction for drug synthesis [64, 65], or a gene annotator identifying the gene for an "orphan" enzyme with known function but unknown sequence [66]. For Task 2, a query reaction is passed through a trained model to perform retrieval to an EC number. The EC number is then associated with proteins that are likely to be able to perform that reaction.
Splits for Task 2. Task 2 aims to evaluate how well a model generalizes to unseen reactions with different levels of difficulty. The train-test splits for Task 2 (easy, medium, and hard) are summarized in Table 3 and visualized in Appendix Figure A.3B. We equate greater difficulty with a more challenging train-test split; in a harder set, the test reactions are less similar to reactions in the corresponding train set. We decide similarity based on the amount of overlap in EC number (e.g., a reaction from 4.2.1.20 is more similar to another one in 4.2.1.20 than one from 4.2.1.1).
@@
Easy Certain reactions are held out, sampled uniformly across ECs, but no EC numbers are held out. The test set is the same as the holdout set. 4,673 58,945 177
Medium All reactions corresponding to certain ECs are held out, at EC level 4 (X.X.X.X). Test set reactions are sampled uniformly across ECs from the holdout set. 4,496 55,851 177
Hard All reactions corresponding to certain ECs are held out, at EC level 3 (X.X.X.-). Test set reactions are sampled uniformly across ECs from the holdout set. 3,326 39,893 163
@@
The sequences in the test splits are distributed generally evenly across different EC numbers (Appendix Figure A.4). From easy to medium to hard, the test set reactions also become more dissimilar to their respective training sets (Figure 2B). Reaction similarity was quantified by DRFP, which is a reaction representation that uses set differences between product and reactant fingerprints and has demonstrated solid performance with requiring model training [62]. Overall, Task 2 is more complex than Task 1, because unlike Task 1, entire EC numbers are held out from the training set (i.e., the classification output vocabulary is open rather than closed), so multiple modalities must be considered. More details on splitting can be found in Appendix A.2.
CREEP baseline method. There are no existing models that have been tested for their ability to generalize beyond annotated reactions, so we develop a contrastive learning method for this task, called Contrastive Reaction-EnzymE Pretraining (CREEP) (Figure 3). Our approach is related to CLIPZyme [55], which uses contrastive alignment of reactions represented as 2D graphs and protein structures represented as 3D graphs. CREEP leverages finetuning of pretraining language models, rxnfp [67] and ProtT5 [44], respectively, to learn aligned representations of reactions and proteins. Rxnfp is a BERT-style [68] language model that was trained on reactions represented as SMILES/SMARTS strings and has demonstrated state-of-the-art performance on reaction type classification. ProtT5 is a T5 protein language model that has been used for prediction of various protein properties. We used these language models for simplicity and their ease of finetuning. Uniquely, CREEP can learn on a third modality (CREEP w/ Text): textual description of the EC number based on gene ontology [69], which is encoded using SciBERT [70]. More details on CREEP training can be found in Appendix A.3.
@@
Task 2 benchmarking results. Benchmarking results for Task 2 are summarized in Table 4. We start with two methods as baselines: randomly guessing a ranking of EC numbers (Random) and finding the most similar reaction in the train set to the test query (Similarity Baseline). For the Similarity Baseline, we used DRFP to represent chemical reactions [62]. Details on the downstream multimodal retrieval process can be found in Appendix A.3 and Figure A.4.
Table 4: Performance of various methods on Task 2. Performance is measured as k=1 retrieval accuracy (%). *denotes that the model was not retrained, so there may be data leakage causing performance to be inflated. Bolded accuracy is the best model without potential data leakage.
@@
Easy Random 0.0 0.0 4.0 23.7
Similarity Baseline 60.5 78.0 87.0 93.2
ChatGPT* 1.7 15.8 38.4 62.7
ChatGPT (w/ Text)* 5.6 44.1 70.6 94.9
CLIPZyme* 27.7 45.2 60.5 80.8
CREEP 48.0 70.6 84.2 93.2
CREEP (w/ Text) 60.5 87.0 93.2 97.7
Medium Random 0.0 0.6 3.4 24.9
Similarity Baseline 0.0 40.1 56.5 75.7
ChatGPT* 0.0 14.1 36.7 64.4
ChatGPT (w/ Text)* 7.9 44.6 73.4 95.5
CLIPZyme* 27.7 45.2 60.5 80.8
CREEP 3.4 39.0 58.8 83.1
CREEP (w/ Text) 5.1 54.2 73.4 91.5
Hard Random 0.0 0.0 1.8 17.2
Similarity Baseline 0.0 0.0 14.7 43.6
ChatGPT* 1.2 16.0 26.4 48.5
ChatGPT (w/ Text)* 7.4 54.6 70.6 92.0
CLIPZyme* 8.0 39.3 48.5 73.0
CREEP 1.8 9.2 27.6 62.0
CREEP (w/ Text) 2.5 12.9 31.9 61.3
@@
Overall, the Similarity Baseline is quite strong and performs much better than Random. ChatGPT performs generally well, but it has access to held-out information, which means that it is not the best measure of generalization to unannotated reactions. ChatGPT was prompted with reactions written using compound IUPAC names rather than SMILES strings, both with (ChatGPT w/ Text) and without (ChatGPT) textual descriptions from gene ontology [69]. We also included CLIPZyme [55] as a baseline, but the performance is likely inflated, as we were not yet able to retrain the model.
@@
On the more difficult test sets, CREEP offers an advantage compared to the Similarity Baseline, and particularly when combined with textual description: CREEP (w/ Text). This suggests that utilizing protein sequence information as a modality is useful. Still, performance on the harder splits is quite weak across the board, which suggests that there is significant room for future improvement. We anticipate that contrastive alignment with textual descriptions will play an increasingly important role in enzyme retrieval [54, 53, 47], and there is opportunity for better curation of these descriptions. Combining other modalities, such as reactions and protein structures as graphs as in CLIPZyme, could also yield improvements in performance. Additional details on the implementation of each method are in Appendix A.4, and additional benchmarking results are presented in Figure A.6.
@@ 6 Conclusion
Predicting the functions of enzymes is important for many applications ranging from gene annotation to enzyme engineering. While many models exist to classify enzyme function via EC numbers, there are no standardized benchmarks for evaluation of these models. Furthermore, no existing models have been tested for generalization beyond annotated reactions. To address this need, we introduce CARE, which is a benchmarking suite to formalize model evaluation for these two tasks. We also present CREEP, a model which uses multimodal contrastive learning and is one of the first models that can perform the latter task. A more detailed discussion of limitations and future work is included in Appendix A.6. Overall, CARE is an important tool for encouraging progress in enzyme functional annotation. We believe that we are just seeing the beginning of the widespread adoption of multimodal models for protein functional prediction, and we expect that many researchers will find CARE useful for formulating and evaluating their models.
@@ A.1 Dataset Processing
Our data processing and splitting workflow is shown in Figure A.1. Protein sequence data, paired to EC number(s), were downloaded from UniProt, selecting only reviewed sequences in Swiss-Prot, resulting in a total of 571,282 sequences on the 13th of May 2024. These were filtered to only retain sequences with an EC number. After, sequences were filtered to length between 100 and 650, partial EC numbers (containing a dash) were removed, and duplicate entries (based on EC and sequence) were dropped, leaving 183,162 sequence-EC pairs. We call this the protein2EC dataset.
@@
Reaction data were downloaded from ECReact and EnzymeMap. ECReact contains 62,222 entries of reaction in SMILES as SMARTS strings format coupled with EC number from a range of data sources. EnzymeMap had 349,458 entries containing sequence-reaction-EC triplets with the reaction in several formats, SMILES as SMARTS string, an atom-level mapping of bond formation and breaking (mapped reaction), and reactions written as text based on IUPAC compound names. Starting with EnzymeMap, protein sequence information was dropped, and duplicate reaction-EC pairs were removed leaving 62,896 reaction-EC pairs. EC numbers that existed in ECReact but not in Enzyme map were added to EnzymeMap providing our reaction dataset of 72,490 reaction-EC pairs. Note the ECReact only contains reactions formatted as SMILES/SMARTS. We call this the reaction2EC dataset.
Finally, we filtered the protein2EC and reaction2EC datasets, such that only EC numbers present in both datasets are retained, leaving 59,122 reaction-EC pairs and 169,865 sequence-EC pairs, with 4,673 unique EC numbers. Clustering using MMseqs2 [71] was performed at 30%, 50%, 70%, and 90% sequence identity and included in the protein2EC dataset for downstream use. The distributions of the curated datasets are shown in Figure A.2.
@@ A.2 Dataset Splitting Details
Using the results from clustering, we generated four test sets, with approximately <30%, 30-50%, 50-70%, and 70-90% sequence identity, respectively, to sequences in the training set. We also only considered ECs at level 4 which had isolated clusters, in other words, an MMseqs2 cluster with only a single member. For every EC level 3 (X.X.X.-), we sampled a random EC at level 4 (X.X.X.X), and then randomly selected a single isolated sequence within that class to add to the test set. This provided balanced test datasets across the functions at EC level 3. For this set, we also enforced that the sequences did not map to multiple EC numbers. Test sets were balanced across functions (e.g. EC number), rather than across their prevalence as protein sequences in nature, based on the idea that most synthesis planners and enzyme engineers would find more value in predictive ability across a broad range of functions.
Next, we looked for "promiscuous" enzymes, which in this work, we define as those annotated with multiple EC numbers. For our promiscuous test set, we only considered using promiscuous enzymes that had a combination of ECs that occurred at least twice. From this set we only took a single sample from each EC combination, and finally randomly selected a single entry from those which have ECs with high surprise levels (i.e. X.X.-.-, where variation occurs at the third EC level or higher).
Finally, we compiled a commonly misclassified test set from Price et al., filtered to sequences of length 100-650, and dropped test sequences that are present in our protein2EC dataset. The pooled set of all proteins in the test sets were removed from our training data leaving 168,682 sequence-EC pairs for training. As a result of the holdout process, 3 of the 4,673 EC numbers are missing from the train set, which should minimally affect performance on the test sets.
@@
For Task 2, we sought to test reaction extrapolation, namely, retrieval of enzymes based on unseen reactions (visualized in Figure A.3B). For the easy and medium test sets we randomly sampled an EC level 4 EC (X.X.X.X) from every level three EC (X.X.X.-). From each of the held-out level 4 ECs, a random reaction was sampled to be in the test set. For the test set, we only considered (1) reactions that do not map to multiple EC numbers, (2) EC numbers with at least two reactions, and (3) reactions from EnzymeMap, which are higher fidelity and consequently have fewer misannotatations.
@@
The hard set was evenly balanced across a random sample from EC level 2 (X.X.-.-) to ensure there was an even distribution across level 3 (EC3, X.X.X.-), from which 54 random EC3s are sampled, with all reactions under the 54 EC3s held out. 171 reactions associated with the 54 ECs were used as the test set, which are evenly spread across the ECs at level 3. Note we ensure the hard test reactions are shared with the easy and medium ones, when possible, to correlate the performance of our test splits. It should be noted that the hard train-test split is different from the easy and medium splits. For Task 2, all of the protein-EC pairs can be used during training.

### v2

@@ 5 Task 2: Enzyme Retrieval
Task 2, enzyme retrieval from a query reaction, tests the ability of a model to extrapolate to unseen reactions. Task 2 has not been formalized or explored in previous studies, but it is equally as important as Task 1, as it applies to a use case where a scientist or engineer is seeking to identify a previously characterized enzyme sequence that can perform a novel (unannotated) reaction (Figure 1D). Typical applications include: an environmental engineer looking for an enzyme to degrade a toxic pollutant [87], an enzyme engineer looking for an enzyme to catalyze a selective reaction for drug synthesis [88, 89], or a gene annotator identifying the gene for an "orphan" enzyme with known function but unknown sequence [90]. For Task 2, a query reaction is passed through a trained model to perform retrieval to an EC number and its associated proteins that are likely to be able to perform that reaction.
Splits for Task 2. Task 2 aims to evaluate how well a model generalizes to unseen reactions with different levels of difficulty. The train-test splits for Task 2 (easy, medium, and hard) are summarized in Table 3 and visualized in Appendix Figure A.3B. We equate greater difficulty with a more challenging train-test split; in a harder set, the test reactions are less similar to reactions in the corresponding train set. We decide similarity based on the amount of overlap in EC number (e.g., a reaction from 4.2.1.20 is considered more similar to another one in 4.2.1.20 than one from 4.2.1.1).
@@
Easy Certain reactions are held out, sampled uniformly across ECs, but no EC numbers are held out. The test set is the same as the holdout set. 4,960 61,373 393
Medium All reactions corresponding to certain ECs are held out, at EC level 4 (X.X.X.X). Test set reactions are sampled uniformly across ECs from the holdout set. 4,748 57,691 393
Hard All reactions corresponding to certain ECs are held out, at EC level 3 (X.X.X.-). Test set reactions are sampled uniformly across ECs from the holdout set. 3,052 35,252 460
@@
The sequences in the test splits are distributed generally evenly across different EC numbers (Appendix Figure A.4). From easy to medium to hard, the test set reactions also become more dissimilar to their respective training sets (Figure 2B). Reaction similarity was quantified by DRFP, which is a reaction representation that uses set differences between product and reactant fingerprints and has demonstrated solid performance without requiring model training [85]. Overall, Task 2 is more complex than Task 1, because unlike Task 1, entire EC numbers are held out from the training set (i.e., the classification output vocabulary is open rather than closed), so multiple modalities must be considered to link the unseen reactions to their respective EC numbers. More details on splitting can be found in Appendix A.2.
CREEP baseline method. There is a lack of models that have been tested for their ability to generalize beyond annotated reactions and none that have considered reaction, text and sequence together, so we develop a contrastive learning method for this task, called Contrastive Reaction-EnzymE Pretraining (CREEP) (Figure 3). Our approach is related to CLIPZyme [78], which uses contrastive alignment of reactions represented as 2D graphs and protein structures represented as 3D graphs. Somewhat differently, CREEP leverages finetuning of pretrained language models, rxnfp [91] and ProtT5 [48] to learn aligned representations of reactions and proteins, respectively. Rxnfp is a BERT-style [92] language model that was trained on reactions represented as SMILES/SMARTS strings and has demonstrated state-of-the-art performance on reaction type classification [91]. ProtT5 is a T5 protein language model that has been used for prediction of various protein properties and demonstrates similar performance to ESM [48, 47]. We used these language models for simplicity and their ease of finetuning. Uniquely, CREEP can learn on a third modality (CREEP w/ Text): textual description of the EC number based on gene ontology [93], which is encoded using SciBERT [94]. The use of text has recently been show to boost function prediction tasks [95]. More details on CREEP training can be found in Appendix A.3.
@@
Task 2 benchmarking results. Benchmarking results for Task 2 are summarized in Table 4. We start with two methods as baselines: randomly guessing a ranking of EC numbers (Random) and finding the most similar reaction in the train set to the test query (Similarity Baseline). For the Similarity Baseline, we used DRFP to represent chemical reactions [85]. Details on the downstream multimodal retrieval process can be found in Appendix A.3 and Figure A.4. Methods like CREEP and CLIPZyme retrieve reference protein sequences based on a query reaction, and those proteins are mapped to a ranking of retrieved EC numbers; thus, they can generalize to EC numbers that are not linked to reactions in (i.e. missing from) the training set.
Table 4: Performance of various methods on Task 2. Performance is measured as k=1 retrieval accuracy (%). *denotes that there may be data leakage causing performance to be inflated. Bolded accuracy is the best model.
@@
Easy Random 0.0 1.0 4.6 22.9
Similarity Baseline 59.3 77.1 85.2 90.6
ChatGPT* 4.8 22.6 43.5 71.0
ChatGPT (w/ Text)* 13.7 56.7 81.4 98.5
CLIPZyme 12.2 39.9 61.8 79.9
CREEP 39.4 66.4 79.9 92.9
CREEP (w/ Text) 60.3 89.3 93.9 96.7
Medium Random 0.0 0.5 4.1 17.0
Similarity Baseline 0.0 40.2 55.7 73.3
CLIPZyme 2.0 26.0 46.6 69.0
CREEP 4.1 44.3 63.1 86.5
CREEP (w/ Text) 7.4 59.5 75.6 92.1
Hard Random 0.0 0.9 1.5 18.3
Similarity Baseline 0.0 0.0 13.5 42.0
CLIPZyme 1.1 4.1 13.5 46.7
CREEP 1.3 4.8 18.7 57.6
CREEP (w/ Text) 1.3 9.8 22.2 57.2
@@
Overall, the Similarity Baseline is quite strong and performs much better than Random. ChatGPT was prompted with reactions written using compound IUPAC names rather than SMILES strings, both with (ChatGPT w/ Text) and without (ChatGPT) textual descriptions from gene ontology [93]. We only show ChatGPT performance for the easy split, as ChatGPT is not a good measure of generalization on the medium and hard splits. These both require "leaving out" entire EC numbers before training to ensure there is no data leakage.
@@
On the more difficult test sets, CREEP offers an advantage compared to the Similarity Baseline, and particularly when combined with textual description: CREEP (w/ Text). This suggests that utilizing protein sequence information as a modality is useful. Still, performance on the harder splits is weak across the board, which suggests that there is significant room for future improvement, although retrieval of EC class may not be the ideal metric for out-of-distribution reaction classification. We anticipate that contrastive alignment with textual descriptions will play an increasingly important role in enzyme retrieval [77, 76, 64], and there is opportunity for better curation of these descriptions. Alternatively, CLIPZyme [78] represents reactions and proteins as graphs, but performance is lower, potentially due to some sample loss from some data being unable to be processed to graphs during training and inference. Here, performance could likely be improved by optimizing training hyperparameters. Additional details on the implementation of each method are in Appendix A.4, and additional benchmarking results are presented in Figure A.6.
6 Discussion
We made an important design choice in the CARE benchmarks: to perform classification and retrieval at the coarse-grained level of EC numbers, which could be a limitation. Enzymes can often perform many reactions, meaning it is an acceptable assumption that enzymes belonging to the same EC number will share the capacity to perform similar reactions, even if they are not directly annotated for all reactions. The ultimate task in this domain is to perform direct reaction to protein sequence retrieval and vice-versa (as in CLIPZyme [78] and Reactzyme [96]). However, currently the data for validation is limited. As experimentalists obtain higher resolution annotations of proteins and their associated reactions, this will become more realistic. Some EC numbers are also incorrectly annotated, which are discussed in more detail here [61].
In the future, the proposed train-test splits could be refined for both tasks. For the Task 1 splits, there is some leakage in the sequence identity, with some sequences in test sets lying outside of the enforced sequence identity ranges, likely due to the different sequence similarity algorithms used (MMseqs2 vs Diamond BLAST). MMseqs2 utilizes cascaded clustering which pre-filters based on initial clusters in the target set [97], while BLAST attempts to identify the closest sequence within the specified set based on local similarity. Another limitation is that while there are no duplicate reactions present in both the train and test sets, some of the test reactions are very similar to reactions in the training set despite having different ECs. It would be beneficial to do a more detailed analysis of reaction similarities and explore other representations of reactions to understand which reactions can be considered equivalent. Many EC classes also involve multi-complex enzymes; in other words, certain subunits of these enzyme are not actually performing catalytic activity. Future work could filter out some of the non catalytically active subunits or act to specifically predict the catalytic subunit, or the entire complex based on a reaction. Over time, the train-test splits should be updated as additional functional annotations are acquired and compiled in databases. For example, while over 36 million sequences in BRENDA/UniProt have EC numbers associated with them, these are often detected using homology based models, which may incorrectly assign EC numbers. A more detailed analysis of the specific failure modes of these ML models could be valuable future work [27].
We opted to assess performance using accuracy due to its simplicity and interpretability, but other retrieval and virtual screening metrics such as BEDROC and enrichment could also be explored. For the promiscuous enzymes with multiple EC numbers, we reported accuracy averaged across all of the true labels, but other classification metrics such as precision and recall should be considered in future work. The number of retrieved ECs could also be chosen using strategies like max separation and p-value as implemented in CLEAN [20].
There is also significant opportunity to use other modalities such as textual description and protein structure for more effective representations of enzymes and reactions, or to improve the text-based annotations. The textual annotations used in this work are direct textual references to the EC class, meaning that the models in task 2 that incorporate text (e.g. CREEP), are effectively learning the relationship between the EC as a word and the EC as a number, in addition to the similarity of the reaction. Other tools such as Pika which use multiple types of textual description extend this to learn relations about the enzyme, and could be used in future iterations of CREEP. Future work will involve incorporating structure and graph based representations [98] into CREEP, similar to those used in CLIPZyme [78]. We also plan to do a more detailed analysis of the representations learned by CREEP. The addition of textual description in CREEP potentially introduces indirect data leakage between the train and test sets, so future iterations of CARE will need to consider this. Future work should also consider how to include protein function prediction models that go beyond enzymes–or models that would like to use additional data/modalities [57, 99]–into the CARE evaluation framework. LAB-Bench provides a framework for evaluating scientific reasoning using language, which will increasingly intersect with CARE [72]. A major bottleneck we encountered was that many models, including language models, were not available or difficult to use, limiting our ability to include them as benchmarks. There is also room to improve the prompt engineering of language models which could be further explored to enhance the performance of these models.
We finally note that methods used to retrieve dangerous proteins that could be used as bioweapons or to synthesize dangerous chemicals is a concern. The implications of this are discussed here [100].
@@ 7 Conclusion
Predicting the functions of enzymes is important for many applications ranging from gene annotation to enzyme engineering. While many models exist to classify enzyme function via EC numbers, there are no standardized benchmarks for evaluation of these models. Furthermore, no existing models have been tested for generalization beyond annotated reactions. To address this need, we introduce CARE, which is a benchmarking suite to formalize model evaluation for these two tasks. We also present CREEP, a model which uses multimodal contrastive learning and is one of the first models that can perform the latter task. We encourage developers to integrate their current and future methods or benchmarking results into the CARE Github repository (https://github.com/jsunn-y/CARE/). Overall, CARE is an important tool for encouraging progress in enzyme functional annotation. We believe that we are just seeing the beginning of the widespread adoption of multimodal models for protein functional prediction, and we expect that many researchers will find CARE useful for formulating and evaluating their models.
@@ A.1 Dataset Processing
Our data processing and splitting workflow is shown in Figure A.1. Protein sequence data, paired to EC number(s), were downloaded from UniProt, selecting only reviewed sequences in Swiss-Prot, resulting in a total of 571,282 sequences on the 13th of May 2024. These were filtered to only retain sequences with an EC number that were not annotated as fragments. After, sequences were filtered to length between 100 and 1024, partial EC numbers (containing a dash) were removed, duplicate entries (based on EC and sequence) were dropped, and only EC numbers with associated reactions were kept, leaving 200,654 sequence-EC pairs. We call this the protein2EC dataset.
@@
Reaction data were downloaded from ECReact and EnzymeMap. ECReact contains 62,222 entries of reaction in SMILES as SMARTS strings format coupled with EC number from a range of data sources. EnzymeMap had 349,458 entries containing sequence-reaction-EC triplets with the reaction in several formats, SMILES as SMARTS string, an atom-level mapping of bond formation and breaking (mapped reaction), and reactions written as text based on IUPAC compound names. Starting with EnzymeMap, protein sequence information was dropped (leaving reaction-EC pairs), and duplicate reaction-EC pairs were removed leaving 62,896 reaction-EC pairs. EC numbers that existed in ECReact but not in Enzyme map were added to EnzymeMap providing our reaction dataset of 72,490 reaction-EC pairs. We did this because EnzymeMap is a higher quality reaction dataset, with fewer incorrect entries. Note the ECReact only contains reactions formatted as SMILES/SMARTS, not as mapped reactions. Together, we call this the reaction2EC dataset.
Finally, we filtered the protein2EC and reaction2EC datasets, such that only EC numbers present in both datasets are retained, leaving 61,766 reaction-EC pairs and 185,995 sequence-EC pairs, with 4,960 unique EC numbers. Clustering using MMseqs2 [97] was performed at 30%, 50%, 70%, and 90% sequence identity and included in the protein2EC dataset for downstream use. The distributions of the curated datasets are shown in Figure A.2.
@@ A.2 Dataset Splitting Details
Using the results from clustering, we generated two test sets, with approximately <30% and 30-50% sequence identity, respectively, to sequences in the training set. We also only considered ECs at level 4 which had isolated clusters, in other words, an MMseqs2 cluster with only a single member. For every EC level 3 (X.X.X.-), we sampled up to three random ECs at level 4 (X.X.X.X), and then randomly selected a single isolated sequence from each class to add to the test set. This provided balanced test datasets across the functions at EC level 3. For this set, we also enforced that the sequences did not map to multiple EC numbers. Test sets were balanced across functions (e.g. EC number), rather than across their prevalence as protein sequences in nature, based on the idea that most synthesis planners and enzyme engineers would find more value in predictive ability across a broad range of functions.
Next, we looked for "promiscuous" enzymes, which in this work, we define as those annotated with multiple EC numbers. For our promiscuous test set, we only considered using promiscuous enzymes that had a combination of ECs that occurred at least twice. From this set we only took a single sample from each EC combination, and finally randomly selected a single entry from those which have ECs with high surprise levels (i.e. X.X.-.-, where variation occurs at the third EC level or higher). From the test sets, we did not include sequences that were labeled as belonging to part of a heterogeneous multi-complex protein.
Finally, we compiled a commonly misclassified test set from Price et al., filtered to sequences of length 100-1024, and dropped test sequences that are present in our protein2EC dataset. The pooled set of all proteins in the test sets were removed from our training data leaving 184,529 sequence-EC pairs for training. As a result of the holdout process, a few of the 4,960 EC numbers are missing from the train set, which should minimally affect performance on the test sets.
@@
For Task 2, we sought to test reaction extrapolation, namely, retrieval of enzymes based on unseen reactions (visualized in Figure A.3B). For the easy and medium test sets, we randomly sampled an EC level 4 EC (X.X.X.X) from every level three EC (X.X.X.-). From each of the held-out level 4 ECs, three random reactions were sampled to be in the test set. For the test set, we only considered (1) reactions that do not map to multiple EC numbers, (2) EC numbers with at least four reactions, and (3) reactions from EnzymeMap, which are higher fidelity and consequently have fewer misannotatations.
@@
The hard set was evenly balanced across a random sample from EC level 2 (X.X.-.-) to ensure there was an even distribution across level 3 (EC3, X.X.X.-), from which 53 random EC3s are sampled, with all reactions under the 53 EC3s held out. 460 reactions associated with the 53 ECs at level 3 were used as the test set, which are evenly spread across the ECs at level 3, with up to three reactions from each EC number at level 4. Note we ensure the hard test reactions are shared with the easy and medium ones, when possible, to correlate the performance of our test splits. It should be noted that the hard train-test split is different from the easy and medium splits. For Task 2, all of the protein-EC pairs can be used during training.